A math quiz has one multiple-choice questions with the answers a,b,c and two true/false questions. What is the probablity of answering all the questions right by guessing?
Answer by sudhanshu_kmr(1152)


total no. of possible ways to answer all the questions = 3 * 2 * 2 = 12
total no. of ways to answer all the question correctly = 1

probability of all the answer are right = 1/12